OpenStudy (anonymous):

A right triangle has legs with 9cm and 40cm what is the length of the hypotenuse?

OpenStudy (anonymous):

i think the answer is 41

OpenStudy (utterly_confuzzled):

You can find the hypotenuse using a^2 + b^2 = c^2 with c being the hypotenuse and a & b being the 2 sides

OpenStudy (utterly_confuzzled):

So 9^2 + 40^2 = c^2 Just solve for c and there's your answer

rishavraj (rishavraj):